OpenStudy (anonymous):

Find an equation for the line. Through (2,6) and perpendicular to y= -5/4x+1 a. -4/5+38/5 b.y=-5/4x+17/2 c.y=5/4x+7/2 d/4/5x+22/5

OpenStudy (anonymous):

¬.¬

OpenStudy (anonymous):

use point slope form of a line.

OpenStudy (anonymous):

Then use the reciprocal of the slope and take the opp sign.

OpenStudy (anonymous):

d??? y=4/5x+22/5

OpenStudy (anonymous):

why do u think d?

OpenStudy (anonymous):

Would it be A. since its the only negative option?

OpenStudy (anonymous):

no...the line is perpendicular.....therefore the signs are opposites

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

you were correct...d

OpenStudy (anonymous):

alright i kinda get it. Its D cause u use the slope formula therefore flips it and it changes the sign.

OpenStudy (anonymous):

find the slope first then find the b of y=mx+b to get an equation of a line

OpenStudy (anonymous):

The slope of a perpendicular line is the reciprocal with the opp sign...

OpenStudy (anonymous):

If this is multiple choice, you can look for the appropriate slope...

OpenStudy (anonymous):

d was the only one with slope of 4/5

OpenStudy (anonymous):

The answer is y=4/5x+22/5. To get the slope of the line you find the opposite reciprocal of the slope of the line given since they are perpendicular. This is 4/5. Use this with the coordinates you have to fill in y=mx+b to find out what b is. When you have this (22/5) you put that, along with the slope (4/5) into the formula again and you'll get y=4/5x+22/5.

OpenStudy (anonymous):

slope intercept form line equation is y = mx +c. m =slope. y= -5/4x+1. Slope m1 = -5/4 the lines are perpendicular so m2 = -1/m1 = 4/5 the line perpendicular to the given line is y =4/5x + c substitute (2, 6) in the equation 6 = 4/5 * 2 + c c = 6 - 8/5 = (30-8)/5 = 22/5 the line equation is y = 4/5x + 22/5 option D is right choice. get free math homework help at http://www.mathskey.com

OpenStudy (anonymous):

Hey naren, where are you from?